HETAV PATEL

Grade 11,

Calculate density of NaCl if edge length of nacl unit cell is 564 pm

Calculate density of NaCl if edge length of nacl unit cell is 564 pm

Grade:12

1 Answers

Arun
5 years ago
The volume of a unit cell, Vcell, is given by: 
Vcell = ZM/dN 
where Z is the formula units per unit cell 
M is the molar mass per formula unit 
d is the density in g cm^-3 
N is the Avogadro constant 

NaCl has a fcc lattice and for fcc, Z = 4 
For NaCl, M = 58.44 g mol^-1 

As the unit cell is a cube, the edge length 564 pm = 3√Vcell 
5.64 x 10^-8 cm = 3√(ZM/dN) 
(5.64 x 10^-8 cm)^3 = (ZM/dN) 
dN = ZM/(5.64 x 10^-8 cm)^3 
d = [ZM/(5.64 x 10^-8 cm)^3]/N 
d = [(4 x 58.44 g mol^-1)/(5.64 x 10^-8 cm)^3)]/6.022 x 10^23 mol^-1 
d = 2.16 g cm^-3